### v1

# Hamiltonian simulation-based quantum-selected configuration interaction for large-scale electronic structure calculations with a quantum computer

Abstract
Quantum-selected configuration interaction (QSCI) is one of the most promising approaches for quantum chemical calculations with the current pre-fault tolerant quantum computers. In the conventional QSCI, the Slater determinants used for the wave function expansion are sampled by iteratively performing approximate wave function preparation and subsequent measurement in the computational basis, and then the subspace Hamiltonian matrix is diagonalized on a classical computer. In this approach, the preparation of a high-quality approximate wave function is necessary to compute total energies accurately. In this work, we propose a Hamiltonian simulation-based QSCI (HSB-QSCI) to avoid this difficulty. In the HSB-QSCI, the Slater determinants are sampled from quantum states generated by the real-time evolution of approximate wave functions. We provide numerical simulations for the spin-singlet ground state and the first excited spin-triplet state of oligoacenes (benzene, naphthalene, and anthracene), phenylene-1,4-dinitrene, and hexa-1,2,3,4,5-pentaene molecules; these results reveal that the HSB-QSCI is applicable not only to molecules where the Hartree–Fock provides a good approximation of the ground state, but also to strongly correlated systems with multiconfigurational characteristics (i.e., the case where preparing a high-quality approximate wave function is hard). We have also numerically confirmed that the HSB-QSCI is robust to approximation errors of the Hamiltonian simulation, such as Trotter errors and the truncation errors of Hamiltonian term by maximum locality in the localized molecular orbital basis. Hardware demonstrations of the HSB-QSCI are also reported for the hexa-1,2,3,4,5-pentaene molecule using 20 qubits IBM superconducting device. The differences between the HSB-QSCI energy and the CAS-CI value are at most 0.15 kcal mol-1, achieving chemical precision.
I Introduction
Among the diverse topics in the field of quantum computing, quantum chemical calculations of atoms and molecules have attracted attention as a promising application of quantum computers. A method to perform the full-configuration interaction (full-CI) calculation using the quantum phase estimation (QPE) algorithm was proposed in 2005 [1], and proof-of-principle experiments for the full-CI/STO-3G of the H2 molecule were reported in 2010 [2, 3]. The quantum circuit for QPE-based full-CI is usually very deep, so highly accurate QPE demonstrations were limited to small models with a few qubits [4, 5, 6, 7, 8]. Recently a QPE-type algorithm has been demonstrated in models with up to 33 qubits [9], however, it is still challenging to compute total energies of larger molecules with chemical precision (an error less than 1 kcal mol-1) on quantum computers available today.
To reduce the computational load of quantum computers, several quantum–classical hybrid algorithms have been proposed. In 2014, the variational quantum eigensolver (VQE) was proposed for quantum chemical calculations using noisy intermediate-scale quantum (NISQ) [10] devices. In VQE, the quantum state corresponding to an approximate wave function is generated using a parameterized quantum circuit, and the energy expectation value is computed by repeatedly performing state preparation and measurements. The parameters are then optimized to minimize the energy expectation value on a classical computer. These steps are iterated until convergence is reached. VQE has been extensively studied both theoretically and experimentally [11], but it often suffers from challenges related to the high sampling cost required to achieve chemical precision [12] and issues with variational optimization in the presence of barren plateaus [13].
As an alternative approach to quantum chemical calculations using current quantum devices, the quantum-selected configuration interaction (QSCI) method has been proposed [14]. In the QSCI, quantum computers are used to sample Slater determinants that are important contributors to the ground state wave function. This can be done by running a quantum circuit to prepare an approximate wave function and measuring the quantum state in the computational basis. The subspace Hamiltonian matrix is then constructed from the Slater determinants selected via quantum computation and diagonalized on a classical computer. In this approach, the quantum computer is anticipated to efficiently sample a polynomial number of important Slater determinants from an exponentially large Hilbert space, and the nature of the quantum states used for sampling determines the accuracy of the calculation. At the stage of preparing an approximate wave function in QSCI, we may employ VQE or adiabatic state preparation method [14]. To advance the former technique, combining an adaptive strategy for ansatz construction (ADAPT [15]) to enhance the efficiency of VQE [16], and using a simpler cost function for VQE parameter optimization [17] have also been proposed. However, these approaches may still face challenges in variational optimizations when implemented on quantum hardware. It should also be noted that in the latter approach [17], the VQE wave function with optimized parameters does not necessarily correspond to the minimum with respect to the given Hamiltonian.
Another strategy to prepare the quantum state for QSCI is to perform the coupled cluster singles and doubles (CCSD) calculation on a classical computer and embed the CCSD wave function using the local unitary cluster Jastrow (LUCJ) ansatz [18]. By using the LUCJ ansatz and introducing an error mitigation technique called as a self-consistent configuration recovery (SCCR), which is inspired by the structure of chemistry problems, Robeldo-Moreno and coworkers reported the QSCI calculations on the IBM superconducting quantum processor and supercomputer “Fugaku” for the nitrogen molecule (N2) and the iron–sulfur clusters [2Fe–2S] and [4Fe–4S] with up to 77 qubits [19]. This approach has also been applied to interaction energy [20] and excited state calculations [21, 22]. The combination of the QSCI with density matrix embedding theory has also been reported [23]. While this approach appears to be successful, it is not clear whether it is superior to the method of sampling important Slater determinants from the CCSD wave function on a classical computer.
In this work, we report a Hamiltonian simulation-based quantum state preparation for QSCI (HSB-QSCI), in which important Slater determinants are sampled from the quantum states after the real-time evolutions of initial wave functions. Our proposed approach has several advantages over existing methods: (1) A simple approximate wave function, such as the HF wave function, can be used as the initial wave function, and no variational optimization in VQE is required. (2) Different samples can be obtained by changing the duration of the evolution time. (3) The accuracy of the Hamiltonian simulation does not need to be very high. As we will show later, the first-order Trotter decomposition with the time length of a single Trotter step is sufficient to collect important Slater determinants for the organic molecules under study. In addition, Hamiltonian term truncation based on operator locality in the localized molecular orbital basis can help to reduce the computational cost. (4) Feedback of the QSCI results into quantum computations is possible; that is, Slater determinants found to be important in the QSCI wave function can be added to the initial wave functions. (5) Although short-time evolution can be accurately simulated on a classical computer, long-time evolution poses significant challenges for classical simulation. Note that methods of sampling basis states for Hamiltonian matrix diagonalization based on Hamiltonian simulation have already been investigated in the framework of quantum Krylov subspace algorithms [24, 25]. Unlike the quantum Krylov method, which computes and on a quantum computer by sampling, the HSB-QSCI method requires only measurements in the computational (Pauli-Z) basis and is thus less computationally demanding. Here, , where , is the initial wave function, and is the Hamiltonian of the system. In the context of Monte Carlo simulation, a quantum dynamics Hamiltonian Monte Carlo (qdHMC) method is also proposed [26], in which the proposal step is performed on a quantum computer by sampling after time evolution.
As a proof-of-concept demonstration of the HSB-QSCI method, we performed numerical simulations for the spin-singlet ground state and the first excited spin-triplet state of oligoacenes [benzene (12 qubits), naphthalene (20 qubits), and anthracene (28 qubits)], phenylene-1,4-dinitrene (20 qubits), and hexa-1,2,3,4,5-pentaene (20 qubits) in both planar and twisted geometries. Our numerical simulations showed that the HSB-QSCI can calculate the total energy with chemical precision by sampling the Slater determinants with a few steps of Hamiltonian simulations. We also report hardware demonstrations of the HSB-QSCI for the lowest spin-singlet and spin-triplet states of hexa-1,2,3,4,5-pentaene using ibm_torino processor, with the aid of matrix product operator (MPO)-based classical compression of the quantum circuit for time evolution. The HSB-QSCI energies calculated on ibm_torino processor agree with the CAS-CI values within 0.15 kcal mol-1 of error. For the calculation of spin triplet states, the accuracy of the HSB-QSCI energy was further improved by adding two important Slater determinants in the HSB-QSCI wave function to the initial wave function of the Hamiltonian simulation.
II Theory
In the ab initio molecular orbital theory, the wave function taking into account the electron correlation is expressed by a linear combination of the Slater determinants as follows:
Here, is the HF determinant and are electronically excited determinants from the HF, and are the corresponding expansion coefficients. In the CI method, the number of Slater determinants included in the wave function expansion increases rapidly with the number of molecular orbitals and electrons, and the excitation order. To calculate the total energy with high accuracy and a low computational cost, methods to perform the CI expansion with selected Slater determinants have been investigated [27, 28, 29, 30]. The QSCI method uses a quantum computer to select important Slater determinants through the approximate wave function preparation and the subsequent computational-basis measurement. In the HSB-QSCI method, important Slater determinants are sampled from the measurement of the quantum state obtained from the time evolution of an approximate wave function.
A schematic view of the HSB-QSCI method is shown in Figure 1. In the HSB-QSCI, the electronic configurations (Slater determinants) used as the basis for the Hamiltonian matrix are sampled from the time-evolved wave functions :
To simulate the real time evolution on a quantum computer, second quantized Hamiltonian given as
is transformed to a qubit Hamiltonian
using a fermion–qubit mapping method. Here, and in Eq. (3) are one- and two-electron integrals, respectively, and and are the creation and annihilation operators, respectively, acting on the -th spin orbital. is a tensor product of Pauli operators referred to as a Pauli string, and is its corresponding coefficient computed from and . In this work, we used the Jordan–Wigner transformation (JWT) [31] as the fermion–qubit mapping, where each qubit stores the occupation number of the corresponding spin orbital: if the spin orbital is occupied, and otherwise. In the JWT, the number of qubits required for wave function mapping is equal to the number of spin orbitals in the active space. The quantum circuit for the time evolution operator can be constructed using conventional approximate approaches such as the Trotter–Suzuki decomposition [32, 33] and qDRIFT [34], or prepared through classical optimization [9] or via variational quantum algorithms [35].
The initial wave function should have an overlap with the target electronic state. The HF wave function is a reasonable choice for in the electronic ground state calculations of typical closed-shell singlet molecules in their equilibrium geometries, but need not be a single Slater determinant. As we will discuss below, using an appropriately prepared multiconfigurational wave function as the initial wave function can reduce the number of sampled Slater determinants while keeping the QSCI energy almost unchanged. By defining the initial wave function as
where is the -th eigenfunction of the Hamiltonian, the quantum state after the time evolution is written as follows:
It is clear that the total evolution time length controls the magnitude of the interferences between the eigenstates, and the measurements of the quantum states in the computational basis with different can yield different Slater determinants. The information of the sampled Slater determinants is transferred to the classical computer, and the Hamiltonian matrix with the selected configurations is constructed and then diagonalized it to obtain the QSCI wave functions and energies.
The effectiveness of sampling in the HSB-QSCI is mainly controlled by the evolution time length , the number of time steps , and the initial wave function . Note that from Eq. (6) we expect that must be set longer for systems with smaller energy gaps with the excited states. In the HSB-QSCI, we can reconstruct using the information of the QSCI wave function by constructing a multiconfigurational wave function consisting of several Slater determinants that have large contributions in the QSCI wave function in the previous step. Performing Hamiltonian simulations with different initial wave functions and merging the measurement results to perform the QSCI is another option. The availability of such feedback from the subspace Hamiltonian diagonalization part on a classical computer to the state preparation part on a quantum computer is one of the important features of the HSB-QSCI method.
Note that the Slater determinants are not always eigenfunctions of the electron spin operator (spin eigenfunctions). In fact, open-shell Slater determinants with spin- unpaired electron(s) are not spin eigenfunctions. Here we have assumed that , where and are the numbers of spin- and spin- electrons, respectively. In order to make the QSCI wave function being spin eigenfunctions, we introduced a symmetry completion step before constructing the Hamiltonian matrix. In the symmetry completion step, open-shell Slater determinants that are missing to span the spin eigenfunction are added. For example, if the bit string corresponding to the Slater determinant is sampled (2, , , and 0 stand for the corresponding molecular orbital is doubly occupied, singly occupied by a spin- electron, singly occupied by a spin- electron, and unoccupied, respectively) but the is not, then we add .
In case when the Hamiltonian simulations are carried out on a current noisy quantum hardware, the measurements of the quantum state give bit strings that correspond to the Slater determinants with different numbers of electrons. The SCCR proposed recently [19] can be adopted to recover the Slater determinants with a correct number of electrons. In the SCCR, if the bit string obtained from the measurement has the wrong number of electrons, bits are flipped to recover the desired number of electrons, and the probability of bit flipping is determined from the occupation number of the molecular orbital calculated from the QSCI wave function of the previous step. This process is iterated until convergence or the iteration step reaches the predefined maximum steps. The SCCR is performed before symmetry completion.
III Computational conditions
To demonstrate the HSB-QSCI, we performed numerical simulations on a classical computer. The target systems are the spin-singlet ground state and the first excited spin-triplet state of oligoacenes (benzene, naphthalene, and anthracene), and phenylene-1,4-dinitrene as a representative system of the open-shell singlet ground state, and hexa-1,2,3,4,5-pentaene, as shown in Figure 2. Geometry optimizations of oligoacenes and hexa-1,2,3,4,5-pentaene in planar geometry were performed at the B3LYP/6-31G* level of theory. The twisted geometry of hexa-1,2,3,4,5-pentaene is generated by rotating one of the CH2 moieties 90 degrees from the equilibrium geometry. The geometry optimization of phenylene-1,4-dinitrene was carried out at the CASSCF(10e,10o)/6-31G* level. Here, (e,o) represents the active space consisting of electrons and molecular orbitals. Cartesian coordinates of all molecules are given in Supporting Information.
Since the computational cost of numerical simulations of quantum circuits scales exponentially with the number of qubits, we adopted the active space approximation. We used the (6e,6o), (10e,10o), and (14e,14o) active spaces for benzene, naphthalene, and anthracene, which consist of the valence and orbitals. The one- and two-electron integrals of oligoacenes are calculated at the RHF/STO-3G level. The active space of phenylene-1,4-dinitrene contains the valence and orbitals and the in-plane 2p orbital of the nitrogen atoms, and the size is (10e,10o). The CASSCF(10e,10o)/6-31G* molecular orbitals are used to construct the second quantized Hamiltonian of phenylene-1,4-dinitrene. In the case of hexa-1,2,3,4,5-pentaene with (10e,10o) active space, the RHF/6-31G* calculation is performed and then the occupied molecular orbitals are localized using the Pipek–Mezey method [36]. For the virtual orbitals, we first formed the singular value decomposition (SVD) quasi-atomic external orbitals using an SVD with respect to the accurate atomic minimal basis functions, and then formed the ordered external orbitals using exchange integrals (using a keyword EXTLOC=ATMNOS in GAMESS-US software [37]). The active orbitals of all molecules are provided in Figures S1–S6 in Supporting Information. One- and two-electron atomic integrals are obtained from GAMESS-US, and the AO–MO integral transformations were performed using our own Python3 program. The qubit Hamiltonian is then generated using the OpenFermion library [38]. All the DFT calculations were done with the Gaussian 16 package [39], and the RHF and CASSCF calculations were carried out with the GAMESS-US software.
As we mentioned in the previous section, the time evolution operation in the HSB-QSCI does not have to be exact, and we can introduce various approximations. In this work, we investigated the effect of Hamiltonian truncation on the sampled Slater determinants in the Hamiltonian simulation and on the QSCI energies. We investigated Hamiltonian truncation based on the locality of the qubit Hamiltonian terms. In this approach, we first generate the localized molecular orbitals (LMOs), and then rearrange the LMOs according to their relative spatial distances. During the qubit Hamiltonian construction using the reordered LMOs, we calculate the locality of the Pauli string (number of Pauli-X, Y, and Z operators in the Pauli string). If the locality of the Pauli string exceeds the threshold, then we exclude it from the qubit Hamiltonian. We investigated this method in hexa-1,2,3,4,5-pentaene, keeping in mind to perform the Hamiltonian simulations on a superconducting quantum device with MPO-based classical compression of the time evolution quantum circuit. The procedure of this step is illustrated in Figure 3.
To sample the Slater determinants based on the real-time evolution of the initial wave function, we set the evolution time for to be and . To construct the quantum circuit for , we adopt the first-order Trotter decomposition with a single Trotter slice, unless otherwise specified. Magnitude ordering [40] is used for the Trotterized term ordering. The quantum circuit simulations were done using the qsim library [41], which allows us to use GPGPU. The number of shots for each is set to be unless otherwise noted. The QSCI part including SCCR, Hamiltonian matrix construction, and diagonalization was done using the qiskit-addon-sqd library [42].
IV Results and Discussion
IV.1 The spin-singlet ground (S0) state and the first excited spin-triplet (T1) state of oligoacenes
There is no doubt that aromatic rings are one of the most important molecular skeletons in chemistry, and the study of the electronic structures of oligoacenes is very important. It is known that the zigzag edge of graphene fragments exhibits strong open-shell characters [43], and the contribution of the HF electronic configuration to the full-CI wave function decreases for larger oligoacenes. Electronic excited states of oligoacenes are also of interest because of their potential for various applications such as singlet fission photovoltaics [44].
The results of the HSB-QSCI simulations of oligoacenes are shown in Figure 4. The RHF and ROHF-like single configurational wave functions are used as the starting wave functions for the Hamiltonian simulations of the S0 and T1 states, respectively. The number of Slater determinants in the CAS-CI wave function is 104 (S0) and 61 (T1) for benzene, 15912 (S0) and 11076 (T1) for naphthalene, and 2945056 (S0) and 2255121 (T1) for anthracene, in the point group. Our numerical simulations revealed that the Slater determinants that are important to describe the target electronic states are efficiently sampled from the Hamiltonian simulations. In benzene, the system size is very small and all possible Slater determinants are sampled up to the 7th steps. The is nonzero even if all Slater determinants are included in the QSCI, due to numerical discrepancy, specifically, the different program used for the CAS-CI and the QSCI. We used the eigsh function in SciPy [45] for the CAS-CI energy calculations, while the QSCI is performed using the selected CI module in PySCF [46]. Importantly, chemical precision () was achieved with three steps of the Hamiltonian simulations in all three molecules studied. In anthracene with three steps of the time evolutions, the HSB-QSCI considers 23607 and 22935 determinants for S0 and T1, respectively, which corresponds to 0.80% and 1.02% of the total determinants. The number of Slater determinants considered in the HSB-QSCI calculations is smaller than that in the CISDTQ calculation within the same active space (91601 and 66852 for S0 and T1 states, respectively), but the HSB-QSCI energy is much lower than the CISDTQ values. The values of the CISDTQ with (14e,14o) active space were calculated to be 5.17 and 18.80 kcal mol-1 for the S0 and T1 states, respectively. Interestingly, we found that up to 7-electron excited determinants were sampled in a Hamiltonian simulation with in anthracene. The ability of Hamiltonian simulations to capture higher-order excitations is astounding. These results illustrate the ability of the Hamiltonian simulation to sample a polynomial number of important Slater determinants from exponentially large Hilbert space. The error in the energy can be systematically reduced by increasing the time steps. We have also studied the dependence of the HSB-QSCI energy on the number of shots in anthracene. By increasing the number of shots to , the number of Slater determinants sampled from the Hamiltonian simulation increases by a factor of about 1.5, and the HSB-QSCI energy decreases systematically (see Figure S7 in Supporting Information).
IV.2 Open-shell spin-singlet and excited spin-triplet states of phenylene-1,4-dinitrene
Phenylene-1,4-dinitrene has two major resonance structures, the diradical with a quinonoid skeleton as described in Figure 2, and the dinitrene with an aromatic ring. Its electronic ground state is an open-shell spin-singlet state with two unpaired electrons in in-plane 2p orbitals of the nitrogen atoms. The thermally excited spin-triplet state of phenylene-1,4-dinitrene has been observed by ESR spectroscopy [47, 48, 49], and the singlet–triplet energy gap is determined to be 0.82 kcal mol-1 from the Curie analysis of the ESR signal [50]. The experimentally determined zero-field splitting parameter in the excited spin-triplet state is = 0.171 cm-1 [47], which is about 10 times larger than the value estimated from a point-dipole approximation. This deviation was explained by the lack of electron correlation effect in the point-dipole approximation [51], insisting that sophisticated consideration of electronic correlation is essential to describe its electronic structure.
In the CASSCF optimized orbital basis, the CAS-CI wave function of the electronic ground state is mainly described by the HF configuration () and the HOMO–LUMO two-electron excited configuration from the HF configuration () (see Table S13 in Supporting Information). In the HSB-QSCI simulations of the open-shell spin-singlet ground state, we performed Hamiltonian simulations with three different initial wave functions, , , and the CAS-CI(2e,2o) wave function , to investigate the dependence of the initial wave function on the energies and the convergence behavior of the HSB-QSCI method. The QSCI simulations were performed under four different conditions: (1) use the results with shots for each time step, (2) use the with shots, (3) use the samples from with shots, and (4) merge the samples from and with shots for each step. For the lowest spin-triplet state we used the ROHF-like single determinant as the initial wave function for the Hamiltonian simulation. The HSB-QSCI results are summarized in Figure 5. Our numerical simulations suggest that almost the same number of Slater determinants are sampled when and are used. Interestingly, the number of Slater determinants sampled from the Hamiltonian simulation is smaller when is used as the initial wave function, while the accuracy of the HSB-QSCI energy remains almost unchanged. This result suggests that the probability of sampling less important Slater determinants is reduced by using more accurate initial wave functions. The accuracy of the HSB-QSCI energies of the spin-singlet ground state and the first excited spin-triplet state is comparable, and the singlet–triplet energy gap is calculated to be 1.07 kcal mol-1, in good agreement with the CAS-CI value (1.06 kcal mol-1) and the experimental one [50].
IV.3 The spin-singlet ground state and the first excited spin-triplet state of hexa-1,2,3,4,5-pentaene
Our last example is the spin-singlet ground state and the first excited spin-triplet state of hexa-1,2,3,4,5-pentaene. The one-dimensional sp-hybridized carbon atom chain molecules are called as carbyne. From the DFT calculations, the carbyne with =CH2 termination is predicted to have a helical conjugation in the non-planar geometries [52, 53]. In the lowest spin-singlet state, the planar structure is the energy minimum, but in the lowest spin-triplet state, the planar geometry is a saddle point and the twisted geometry becomes stable. As a result, the singlet–triplet energy gap strongly depends on the dihedral angle between two =CH2 terminations. In this study, we calculated the lowest spin-singlet and the first excited spin-triplet states of the geometries with the dihedral angles of and .
Because carbyne is a one-dimensional molecule, using LMOs and adopting Hamiltonian truncation by maximum locality may be a good option to reduce the computational cost of Hamiltonian simulation. In this work, we investigated Hamiltonian term truncation by operator locality of the Pauli string in the qubit Hamiltonian. Since LMOs are reordered by relative distances before constructing a qubit Hamiltonian, operator locality-based truncation is roughly equivalent to Hamiltonian term truncation based on spatial distances. To assess the effect of Hamiltonian truncation with the locality, we first calculated the number of Hamiltonian terms in the second quantized Hamiltonian and the fidelity of the ground state wave function, , where and are the ground state wave functions with the truncated and untruncated Hamiltonians, respectively. The results are summarized in Tables 1 and 2 for planar and twisted geometries, respectively.
| Maximum locality | #Terms | Fidelity |
| 20 | 20072 | 1.0 |
| 18 | 19676 | 0.999816 |
| 16 | 18596 | 0.998764 |
| 14 | 17592 | 0.998571 |
| 12 | 16248 | 0.996394 |
| 10 | 12708 | 0.992327 |
| 8 | 7836 | 0.946569 |
| 6 | 4840 | 0.941506 |
| 4 | 3048 | 0.912734 |
| Maximum locality | #Terms | Fidelity |
| 20 | 40200 | 1.0 |
| 18 | 39772 | 0.999576 |
| 16 | 38244 | 0.997395 |
| 14 | 35232 | 0.996860 |
| 12 | 30640 | 0.995021 |
| 10 | 24660 | 0.987597 |
| 8 | 17772 | 0.898829 |
| 6 | 10744 | 0.891525 |
| 4 | 4632 | 0.843284 |
Note that the number of Hamiltonian terms is larger in the twisted geometry than in the planar one, which is due to the difference in the spatial point group. The planar geometry belongs to , which has eight irreducible representations, while the twisted geometry is calculated in point group with four irreducible representations. As can be seen from Tables 1 and 2, the number of Hamiltonian terms is reduced by about 40% when considering up to 10-local terms, with the fidelity of the ground state wave function being about 0.99. The rapid decrease of the fidelity for the maximum locality smaller than 10 can be explained by the fact that the – and – interactions (here we assumed that one-dimensional carbon atom chain is parallel to the z-axis) with the next nearest neighbor C=C bond (for example, – in Figure 3) are described as 10-local operators in the qubit Hamiltonian.
The results of the HSB-QSCI simulations are shown in Figure 6. In the first excited spin-triplet state calculations, we set the starting wave function to carry two unpaired electrons in the central bonds ( and in Figure 3). It is clear that the number of Slater determinants sampled from the Hamiltonian simulations and the convergence behavior of the HSB-QSCI energies are almost the same for , where is the maximum locality of the qubit Hamiltonian terms. These results are consistent with the trend of the singlet ground state fidelity value. The singlet–triplet energy gap calculated using the real-time evolution with the untruncated Hamiltonian () is 53.93 and 20.26 kcal mol-1 for planar and twisted geometries, respectively, and the difference from the CAS-CI singlet–triplet energy gap is less than 0.01 kcal mol-1. When the maximum locality is set to be = 6 or 8, the error in the HSB-QSCI energy becomes larger, but the HSB-QSCI still achieves chemical precision by considering 10 time steps. The calculated singlet–triplet energy gaps are 54.16 (, planar), 54.04 (, planar), 20.38 (, twisted), and 20.26 (, twisted) kcal mol-1. As a result of partial cancellation of the error in the total energies of the singlet and triplet states, the error in the singlet–triplet energy gap is at most 0.23 kcal mol-1.
It should be emphasized that the ROHF-like single configurational wave function in the LMO basis used in the first excited spin-triplet state calculation has rather a small overlap with the CAS-CI wave function. The squared overlap values in the first excited spin-triplet state are 0.1691 and 0.1240 for planar and twisted geometries, respectively (see Tables S16 and S18 in Supporting Information for the CAS-CI wave function). The fact that the HSB-QSCI provides accurate energy with such a small overlap is promising, because the overlap between the HF and the full-CI wave functions decreases with system size, and the preparation of a sophisticated approximate wave function becomes challenging for larger molecules.
IV.4 Hardware demonstration of the spin-singlet ground state and the first excited spin-triplet state calculations of hexa-1,2,3,4,5-pentaene
Since the Hamiltonian term truncation based on the locality of the qubit Hamiltonian terms works excellently in hexa-1,2,3,4,5-pentaene, we performed the HSB-QSCI calculation on the IBM superconducting quantum processor. To reduce the quantum circuit depth of the Hamiltonian simulation, we performed the MPO-based classical compression of the quantum circuit. This approach was demonstrated in our previous study on the quantum phase difference estimation-based energy gap calculations [9]. The details of the MPO-based quantum circuit optimization are given in the reference [9]. In this work, we used a Hamiltonian truncated by the maximum locality ( = 6), and the 6-layer brick wall type quantum circuit is generated for the time evolution operator of . The number of sweeps on the optimizations was set to 10000. In constructing the MPOs, the second-order Trotter decomposition with a single Trotter slice was employed, and the singular value cutoff for the SVD was set to . The order of terms in the Trotter decomposition is set as follows. First, we generate the truncated second quantized Hamiltonian by checking the one- and two-electron integrals in the lexicographic order: after in , and , , , , in that order in . Then the truncated second quantized Hamiltonian is transformed into a qubit Hamiltonian while retaining the order of the terms, and the Trotterized time evolution operator is constructed subsequently. The number of shots was set to be , and the HSB-QSCI was performed five times with and without the SCCR procedure. We also ran the HSB-QSCI calculations with shots to see the shot number dependence.
We first checked the Hamming weight distribution of the bit strings obtained in the measurements and compared it to the uniform distribution. The Hamming weight distributions at are plotted in Figure 7. We confirmed that the Hamming weight distributions of the bit strings obtained from the Hamiltonian simulation with are significantly different from the uniform distribution. However, due to hardware noise, only 19–26% of the measured bit strings have correct Hamming weights (10 in this case).
The HSB-QSCI results are summarized in Figure 8. The differences of the HSB-QSCI energies without SCCR using shots for each time step from the CAS-CI values are about an order of magnitude larger than the statevector results. However, the error in the energy is systematically reduced by using the SCCR. The error of the HSB-QSCI energies is larger for the spin-triplet state than for the spin-singlet ground state, presumably due to the strong multiconfigurational character in the spin-triplet state with the LMO basis. The accuracy of the HSB-QSCI energies improved as the number of shots increased. In fact, when we performed shots for each time step, the values are calculated to be 0.0046, 0.1482, 0.0015, and 0.0406 kcal mol-1 for planar (spin-singlet), planar (spin-triplet), twisted (spin-singlet), and twisted (spin-triplet), respectively, achieving chemical precision.
Since the lowest spin-triplet state of hexa-1,2,3,4,5-pentane shows a strong multiconfigurational character in the LMO basis, we investigated the feedback from the HSB-QSCI results to improve the initial wave function. Using the QSCI wave function calculated from the real-time evolution data up to , we selected three Slater determinants with the CI coefficient and constructed the three-configurational wave functions, by maintaining the ratio of the coefficients. The resulting wave functions are and . Here, we used the orbital ordering in Figures S5 and S6 for planar and twisted geometries, respectively, to represent the wave function. Since the preparation of requires non-nearest neighbor two-qubit gates, the depth of the quantum circuit increased from 22 to 102, in the twisted geometry with . By using the dynamical decoupling [54] in the quantum circuit execution, the Hamming weight distribution after ten steps of time evolution shows a clear difference from the uniform distribution (see Figure S8 in Supporting Information). Interestingly, the HSB-QSCI with SCCR gave very accurate energy in at most 3 steps (see Figure S9 in Supporting Information). These results strongly support the usefulness of the HSB-QSCI for accurate quantum chemical calculations on a quantum computer.
V Conclusion
In this study, we proposed an HSB-QSCI, which uses a quantum computer to perform Hamiltonian simulation and subsequent measurement of the quantum state in the computational basis to sample the important Slater determinants used to span the subspace Hamiltonian to be diagonalized on a classical computer. Compared to the reported state preparation approaches for the QSCI, the HSB-QSCI is free from variational optimizations of the quantum states, and it works with the simple initial wave function like the HF. Since long-time evolution is in general difficult to simulate on a classical computer, we expect this approach to become powerful as the system size increases. Proof-of-concept simulations were performed for the lowest spin-singlet and spin-triplet states of organic molecules including benzene, naphthalene, anthracene, phenylene-1,4-dinitrene, and hexa-1,2,3,4,5-pentaene, and we achieved chemical precision in all molecules studied. The quantum hardware demonstrations of the HSB-QSCI for the lowest spin-singlet and spin-triplet states of hexa-1,2,3,4,5-pentaene were also reported using ibm_torino processor with 20 qubits, with the aid of MPO-based classical optimization of the quantum circuit for the time evolution operator. By setting the number of shots to and using the SCCR, we obtained the HSB-QSCI total energies with a deviation from the CAS-CI values of at most 0.15 kcal mol-1, illustrating the usefulness of the HSB-QSCI approach.
VI Acknowledgments
This work was supported by Quantum-LEAP Flagship Program (Grant No. JPMXS0120319794 and JPMXS0118067285) from Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan. K.S. acknowledges support from Center of Innovations for Sustainable Quantum AI (JPMJPF2221) from Japan Science and Technology Agency (JST), Japan and Grants-in-Aid for Scientific Research C (21K03407) and for Transformative Research Area B (23H03819) from Japan Society for the Promotion of Science (JSPS), Japan. A part of this work was performed for Council for Science, Technology and Innovation (CSTI), Cross-ministerial Strategic Innovation Promotion Program (SIP), “Promoting the application of advanced quantum technology platforms to social issues” (Funding agency: QST). The authors thank Takashi Abe, Yoshiharu Mori, and Hajime Nakamura for useful discussions. The part of calculations was performed on the Mitsubishi Chemical Corporation (MCC) high-performance computer (HPC) system “NAYUTA”, where “NAYUTA” is a nickname for MCC HPC and is not a product or service name of MCC. We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team.
S1 Supporting Information
S1.1 Active spaces
S1.2 Cartesian coordinates
| Atom | X | Y | Z |
| C | 0.000000 | 1.396602 | 0.000000 |
| C | 0.000000 | 1.396602 | 0.000000 |
| C | 1.209493 | 0.698301 | 0.000000 |
| C | 1.209493 | 0.698301 | 0.000000 |
| C | 1.209493 | 0.698301 | 0.000000 |
| C | 1.209493 | 0.698301 | 0.000000 |
| H | 0.000000 | 2.483720 | 0.000000 |
| H | 0.000000 | 2.483720 | 0.000000 |
| H | 2.150965 | 1.241860 | 0.000000 |
| H | 2.150965 | 1.241860 | 0.000000 |
| H | 2.150965 | 1.241860 | 0.000000 |
| H | 2.150965 | 1.241860 | 0.000000 |
| Atom | X | Y | Z |
| C | 0.000000 | 0.000000 | 0.717063 |
| C | 0.000000 | 0.000000 | 0.717063 |
| C | 0.000000 | 1.244791 | 1.402547 |
| C | 0.000000 | 1.244791 | 1.402547 |
| C | 0.000000 | 1.244791 | 1.402547 |
| C | 0.000000 | 1.244791 | 1.402547 |
| C | 0.000000 | 2.433546 | 0.708438 |
| C | 0.000000 | 2.433546 | 0.708438 |
| C | 0.000000 | 2.433546 | 0.708438 |
| C | 0.000000 | 2.433546 | 0.708438 |
| H | 0.000000 | 1.242227 | 2.490221 |
| H | 0.000000 | 1.242227 | 2.490221 |
| H | 0.000000 | 1.242227 | 2.490221 |
| H | 0.000000 | 1.242227 | 2.490221 |
| H | 0.000000 | 3.378258 | 1.245514 |
| H | 0.000000 | 3.378258 | 1.245514 |
| H | 0.000000 | 3.378258 | 1.245514 |
| H | 0.000000 | 3.378258 | 1.245514 |
| Atom | X | Y | Z |
| C | 0.000000 | 0.000000 | 1.403608 |
| C | 0.000000 | 0.000000 | 1.403608 |
| C | 0.000000 | 1.223844 | 0.722822 |
| C | 0.000000 | 1.223844 | 0.722822 |
| C | 0.000000 | 1.223844 | 0.722822 |
| C | 0.000000 | 1.223844 | 0.722822 |
| C | 0.000000 | 2.479411 | 1.407056 |
| C | 0.000000 | 2.479411 | 1.407056 |
| C | 0.000000 | 2.479411 | 1.407056 |
| C | 0.000000 | 2.479411 | 1.407056 |
| C | 0.000000 | 3.660505 | 0.713049 |
| C | 0.000000 | 3.660505 | 0.713049 |
| C | 0.000000 | 3.660505 | 0.713049 |
| C | 0.000000 | 3.660505 | 0.713049 |
| H | 0.000000 | 0.000000 | 2.491948 |
| H | 0.000000 | 0.000000 | 2.491948 |
| H | 0.000000 | 2.477069 | 2.494589 |
| H | 0.000000 | 2.477069 | 2.494589 |
| H | 0.000000 | 2.477069 | 2.494589 |
| H | 0.000000 | 2.477069 | 2.494589 |
| H | 0.000000 | 4.607098 | 1.246673 |
| H | 0.000000 | 4.607098 | 1.246673 |
| H | 0.000000 | 4.607098 | 1.246673 |
| H | 0.000000 | 4.607098 | 1.246673 |
| Atom | X | Y | Z |
| C | 0.000000 | 0.000000 | 1.429275 |
| C | 0.000000 | 0.000000 | 1.429275 |
| C | 0.000000 | 1.259487 | 0.674849 |
| C | 0.000000 | 1.259487 | 0.674849 |
| C | 0.000000 | 1.259487 | 0.674849 |
| C | 0.000000 | 1.259487 | 0.674849 |
| N | 0.000000 | 0.000000 | 2.710369 |
| N | 0.000000 | 0.000000 | 2.710369 |
| H | 0.000000 | 2.176735 | 1.233246 |
| H | 0.000000 | 2.176735 | 1.233246 |
| H | 0.000000 | 2.176735 | 1.233246 |
| H | 0.000000 | 2.176735 | 1.233246 |
| Atom | X | Y | Z |
| C | 0.000000 | 0.000000 | 3.236615 |
| C | 0.000000 | 0.000000 | 1.918246 |
| C | 0.000000 | 0.000000 | 0.645266 |
| C | 0.000000 | 0.000000 | 0.645266 |
| C | 0.000000 | 0.000000 | 1.918246 |
| C | 0.000000 | 0.000000 | 3.236615 |
| H | 0.000000 | 0.927096 | 3.807401 |
| H | 0.000000 | 0.927096 | 3.807401 |
| H | 0.000000 | 0.927096 | 3.807401 |
| H | 0.000000 | 0.927096 | 3.807401 |
| Atom | X | Y | Z |
| C | 0.000000 | 0.000000 | 3.236615 |
| C | 0.000000 | 0.000000 | 1.918246 |
| C | 0.000000 | 0.000000 | 0.645266 |
| C | 0.000000 | 0.000000 | 0.645266 |
| C | 0.000000 | 0.000000 | 1.918246 |
| C | 0.000000 | 0.000000 | 3.236615 |
| H | 0.000000 | 0.927096 | 3.807401 |
| H | 0.927096 | 0.000000 | 3.807401 |
| H | 0.000000 | 0.927096 | 3.807401 |
| H | 0.927096 | 0.000000 | 3.807401 |
S1.3 The CAS-CI wave function
| Alpha | Beta | Coefficients |
| 111000 | 111000 | 0.9166 |
| 110100 | 110100 | 0.1733 |
| 101010 | 101010 | 0.1733 |
| 110010 | 101100 | 0.1153 |
| 101100 | 110010 | 0.1153 |
| 111000 | 100110 | 0.0829 |
| 100110 | 111000 | 0.0829 |
| 110001 | 011100 | 0.0776 |
| 011100 | 110001 | 0.0776 |
| 101001 | 011010 | 0.0776 |
| 011010 | 101001 | 0.0776 |
| 101010 | 011001 | 0.0558 |
| 011001 | 101010 | 0.0558 |
| 110100 | 011001 | 0.0558 |
| 011001 | 110100 | 0.0558 |
| Alpha | Beta | Coefficients |
| 111100 | 110000 | 0.6456 |
| 111010 | 101000 | 0.6456 |
| 111001 | 011000 | 0.2225 |
| 101110 | 100010 | 0.1341 |
| 110110 | 100100 | 0.1341 |
| 011110 | 100001 | 0.0872 |
| 110101 | 010100 | 0.0811 |
| 101011 | 001010 | 0.0811 |
| 101101 | 010010 | 0.0795 |
| 110011 | 001100 | 0.0795 |
| 111100 | 010100 | 0.0577 |
| 111010 | 010010 | 0.0577 |
| 111100 | 001010 | 0.0577 |
| 111010 | 001100 | 0.0577 |
| 110101 | 110000 | 0.0573 |
| 101101 | 101000 | 0.0573 |
| 101011 | 110000 | 0.0573 |
| 110011 | 101000 | 0.0573 |
| Alpha | Beta | Coefficients |
| 1111100000 | 1111100000 | 0.8615 |
| 1111010000 | 1111010000 | 0.1498 |
| 1110101000 | 1110101000 | 0.1338 |
| 1111001000 | 1110110000 | 0.0973 |
| 1110110000 | 1111001000 | 0.0973 |
| 1111000100 | 1101110000 | 0.0856 |
| 1101110000 | 1111000100 | 0.0856 |
| 1111010000 | 1101100100 | 0.0780 |
| 1101100100 | 1111010000 | 0.0780 |
| 1111100000 | 1110011000 | 0.0735 |
| 1110011000 | 1111100000 | 0.0735 |
| 1110100100 | 1101101000 | 0.0706 |
| 1101101000 | 1110100100 | 0.0706 |
| 1111000010 | 1011110000 | 0.0692 |
| 1011110000 | 1111000010 | 0.0692 |
| 1101100100 | 1101100100 | 0.0671 |
| 1101110000 | 1101110000 | 0.0652 |
| 1111000100 | 1111000100 | 0.0625 |
| 1110100010 | 1011101000 | 0.0587 |
| 1011101000 | 1110100010 | 0.0587 |
| 1110100001 | 0111101000 | 0.0501 |
| 0111101000 | 1110100001 | 0.0501 |
| Alpha | Beta | Coefficients |
| 1111110000 | 1111000000 | 0.7111 |
| 1111101000 | 1110100000 | 0.3703 |
| 1111100100 | 1101100000 | 0.2860 |
| 1111100010 | 1011100000 | 0.1859 |
| 1111100001 | 0111100000 | 0.1176 |
| 1110111000 | 1110001000 | 0.1101 |
| 1111010100 | 1101010000 | 0.1016 |
| 1111011000 | 1110010000 | 0.0952 |
| 1110110100 | 1101001000 | 0.0840 |
| 1101110100 | 1101000100 | 0.0770 |
| 1111010010 | 1011010000 | 0.0715 |
| 1101111000 | 1110000100 | 0.0690 |
| 1110110010 | 1011001000 | 0.0668 |
| 1011111000 | 1110000010 | 0.0637 |
| 1111110000 | 1101001000 | 0.0599 |
| 1110110100 | 1111000000 | 0.0583 |
| 1110101100 | 1100101000 | 0.0556 |
| 1111110000 | 1011010000 | 0.0555 |
| 1111001100 | 1100110000 | 0.0551 |
| 1111010010 | 1111000000 | 0.0542 |
| 1110101010 | 1010101000 | 0.0535 |
| 1111101000 | 1011001000 | 0.0511 |
| 1110110010 | 1110100000 | 0.0505 |
| 1011110010 | 1011000010 | 0.0503 |
| Alpha | Beta | Coefficients |
| 11111110000000 | 11111110000000 | 0.8021 |
| 11111101000000 | 11111101000000 | 0.1588 |
| 11111010100000 | 11111010100000 | 0.1066 |
| 11111100100000 | 11111011000000 | 0.0891 |
| 11111011000000 | 11111100100000 | 0.0891 |
| 11110110010000 | 11110110010000 | 0.0741 |
| 11111100010000 | 11110111000000 | 0.0719 |
| 11110111000000 | 11111100010000 | 0.0719 |
| 11111110000000 | 11111001100000 | 0.0681 |
| 11111001100000 | 11111110000000 | 0.0681 |
| 11111100001000 | 11101111000000 | 0.0673 |
| 11101111000000 | 11111100001000 | 0.0673 |
| 11111100000100 | 11011111000000 | 0.0627 |
| 11011111000000 | 11111100000100 | 0.0627 |
| 11111101000000 | 11110110010000 | 0.0617 |
| 11110110010000 | 11111101000000 | 0.0617 |
| 11111101000000 | 11101110001000 | 0.0579 |
| 11101110001000 | 11111101000000 | 0.0579 |
| 11111010010000 | 11110110100000 | 0.0552 |
| 11110110100000 | 11111010010000 | 0.0552 |
| 11111010001000 | 11101110100000 | 0.0514 |
| 11101110100000 | 11111010001000 | 0.0514 |
| Alpha | Beta | Coefficients |
| 11111111000000 | 11111100000000 | 0.7043 |
| 11111110010000 | 11110110000000 | 0.2348 |
| 11111110100000 | 11111010000000 | 0.2284 |
| 11111110001000 | 11101110000000 | 0.1778 |
| 11111110000100 | 11011110000000 | 0.1439 |
| 11111011100000 | 11111000100000 | 0.0886 |
| 11111110000010 | 10111110000000 | 0.0886 |
| 11111101010000 | 11110101000000 | 0.0865 |
| 11111110000001 | 01111110000000 | 0.0771 |
| 11110111010000 | 11110100010000 | 0.0767 |
| 11111101100000 | 11111001000000 | 0.0749 |
| 11111011010000 | 11110100100000 | 0.0722 |
| 11111101001000 | 11101101000000 | 0.0688 |
| 11111101000100 | 11011101000000 | 0.0634 |
| 11111111000000 | 11011101000000 | 0.0598 |
| 11111101000100 | 11111100000000 | 0.0584 |
| 11111111000000 | 11110100100000 | 0.0581 |
| 11111110001000 | 11111100000000 | 0.0581 |
| 11111111000000 | 11101110000000 | 0.0581 |
| 11111011000100 | 11011100100000 | 0.0571 |
| 11111011010000 | 11111100000000 | 0.0562 |
| 11110111100000 | 11111000010000 | 0.0542 |
| 11111011001000 | 11101100100000 | 0.0533 |
| 11011111100000 | 11111000000100 | 0.0503 |
| Alpha | Beta | Coefficients |
| 1111100000 | 1111100000 | 0.6291 |
| 1111010000 | 1111010000 | 0.6134 |
| 1110101000 | 1110101000 | 0.1358 |
| 1110011000 | 1110011000 | 0.1346 |
| 1111001000 | 1110110000 | 0.0926 |
| 1110110000 | 1111001000 | 0.0926 |
| 1101100100 | 1101100100 | 0.0788 |
| 1101010100 | 1101010100 | 0.0769 |
| 1110100010 | 1011101000 | 0.0666 |
| 1011101000 | 1110100010 | 0.0666 |
| 1110010010 | 1011011000 | 0.0653 |
| 1011011000 | 1110010010 | 0.0653 |
| 1011100010 | 1011100010 | 0.0542 |
| 1011010010 | 1011010010 | 0.0531 |
| 1101100100 | 1110101000 | 0.0513 |
| 1110101000 | 1101100100 | 0.0513 |
| 1110100100 | 1101101000 | 0.0510 |
| 1101101000 | 1110100100 | 0.0510 |
| 1110011000 | 1101010100 | 0.0504 |
| 1101010100 | 1110011000 | 0.0504 |
| Alpha | Beta | Coefficients |
| 1111110000 | 1111000000 | 0.8848 |
| 1110111000 | 1110001000 | 0.1835 |
| 1101110100 | 1101000100 | 0.1105 |
| 1011111000 | 1110000010 | 0.0967 |
| 1110110010 | 1011001000 | 0.0891 |
| 1111101000 | 1110100000 | 0.0840 |
| 1111011000 | 1110010000 | 0.0824 |
| 1101111000 | 1110000100 | 0.0758 |
| 1011110010 | 1011000010 | 0.0742 |
| 1101110100 | 1110001000 | 0.0718 |
| 1110111000 | 1101000100 | 0.0718 |
| 1110110100 | 1101001000 | 0.0671 |
| 1011111000 | 1011001000 | 0.0670 |
| 1011110010 | 1110001000 | 0.0644 |
| 1110111000 | 1011000010 | 0.0633 |
| 0111111000 | 1110000001 | 0.0554 |
| 1101111000 | 1101001000 | 0.0525 |
| 1110110001 | 0111001000 | 0.0500 |
| Alpha | Beta | Coefficients |
| 1111100000 | 1111100000 | 0.8943 |
| 1111000001 | 1111000001 | 0.1638 |
| 0111110000 | 0111110000 | 0.1638 |
| 1101100100 | 1101100100 | 0.1349 |
| 1011101000 | 1011101000 | 0.1266 |
| 1110100010 | 1110100010 | 0.1266 |
| 1101100001 | 1111000100 | 0.0503 |
| 1111000100 | 1101100001 | 0.0503 |
| 0111100100 | 1101110000 | 0.0503 |
| 1101110000 | 0111100100 | 0.0503 |
| Alpha | Beta | Coefficients |
| 1111100100 | 1101100000 | 0.4112 |
| 1111110000 | 0111100000 | 0.3648 |
| 1111100001 | 1111000000 | 0.3648 |
| 1111100100 | 1111000000 | 0.2693 |
| 1111100100 | 0111100000 | 0.2693 |
| 1111110000 | 1101100000 | 0.2677 |
| 1111100001 | 1101100000 | 0.2677 |
| 1111110000 | 1111000000 | 0.1912 |
| 1111100001 | 0111100000 | 0.1912 |
| 1111100100 | 1101000001 | 0.0651 |
| 1111100100 | 0101110000 | 0.0651 |
| 1111000101 | 1101000001 | 0.0643 |
| 0111110100 | 0101110000 | 0.0643 |
| 1111100001 | 1101000100 | 0.0635 |
| 1111110000 | 0101100100 | 0.0635 |
| 1111010001 | 0111000001 | 0.0621 |
| 0111110001 | 0111010000 | 0.0621 |
| 1011101100 | 1001101000 | 0.0585 |
| 1110100110 | 1100100010 | 0.0585 |
| 1101100101 | 1111000000 | 0.0569 |
| 1101110100 | 0111100000 | 0.0569 |
| 1111000101 | 1101100000 | 0.0554 |
| 0111110100 | 1101100000 | 0.0554 |
| 1111110000 | 0111000100 | 0.0528 |
| 1111100001 | 0111000100 | 0.0528 |
| 1111100001 | 0111010000 | 0.0518 |
| 1111110000 | 0111000001 | 0.0518 |
| 1110110010 | 0110100010 | 0.0512 |
| 1011101001 | 1011001000 | 0.0512 |
| 1111100001 | 1101000001 | 0.0508 |
| 1111110000 | 0101110000 | 0.0508 |
| 1011111000 | 0011101000 | 0.0503 |
| 1110100011 | 1110000010 | 0.0503 |
| Alpha | Beta | Coefficients |
| 1111100000 | 1111100000 | 0.8519 |
| 1111000001 | 1111000001 | 0.1630 |
| 0111110000 | 0111110000 | 0.1630 |
| 1101100100 | 1101100100 | 0.1431 |
| 1011101000 | 1011101000 | 0.1295 |
| 1110100010 | 1110100010 | 0.1295 |
| 1101110000 | 0111100100 | 0.0618 |
| 0111100100 | 1101110000 | 0.0618 |
| 1111000100 | 1101100001 | 0.0618 |
| 1101100001 | 1111000100 | 0.0618 |
| Alpha | Beta | Coefficients |
| 1111100100 | 1101100000 | 0.3522 |
| 1111100001 | 1111000000 | 0.2650 |
| 1111110000 | 0111100000 | 0.2650 |
| 1111100100 | 1111000000 | 0.2209 |
| 1111100100 | 0111100000 | 0.2209 |
| 1111100001 | 1101100000 | 0.2074 |
| 1111110000 | 1101100000 | 0.2074 |
| 1111100100 | 1011100000 | 0.1938 |
| 1111100100 | 1110100000 | 0.1938 |
| 1111100010 | 1110100000 | 0.1625 |
| 1111101000 | 1011100000 | 0.1625 |
| 1111110000 | 1011100000 | 0.1498 |
| 1111100001 | 1110100000 | 0.1498 |
| 1111100010 | 1101100000 | 0.1471 |
| 1111101000 | 1101100000 | 0.1471 |
| 1111100001 | 0111100000 | 0.1437 |
| 1111110000 | 1111000000 | 0.1437 |
| 1111100001 | 1011100000 | 0.1228 |
| 1111110000 | 1110100000 | 0.1228 |
| 1111100010 | 1111000000 | 0.1221 |
| 1111101000 | 0111100000 | 0.1221 |
| 1111100010 | 1011100000 | 0.1088 |
| 1111101000 | 1110100000 | 0.1088 |
| 1111100010 | 0111100000 | 0.0962 |
| 1111101000 | 1111000000 | 0.0962 |
| 1111000101 | 1101000001 | 0.0562 |
| 0111110100 | 0101110000 | 0.0562 |
| 1011101100 | 1001101000 | 0.0513 |
| 1110100110 | 1100100010 | 0.0513 |
| 1111100100 | 0101110000 | 0.0511 |
| 1111100100 | 1101000001 | 0.0511 |
S1.4 Shot number dependence on the HSB-QSCI simulations of anthracene
S1.5 Hardware results of the HSB-QSCI with three-configurational wave functions